OpenStudy (devincoolage1001):

-24x+4y+20=0 Solve for y. Please explain.

OpenStudy (kittiwitti1):

"solve for y" = get y by itself, isolate it

OpenStudy (devincoolage1001):

What should the first step to this be?

OpenStudy (kittiwitti1):

\[-24x+4y+20=0\]\[-24x+20+4y-4y=0-4y\]\[-24x+20\cancel{+4y}\cancel{-4y}=\cancel{0}-4y\rightarrow-24x+20=-4y\]

OpenStudy (robtobey2):

Divide each side by 4. -6x+y+5=0 Move the none y terms to the RHS. y = 6x-5

OpenStudy (devincoolage1001):

Thank you both. This was very helpful.